ABC employees are eligible with benefits that provide 15 years’ reimbursement for health care services of $20,000 annually, beginning one year from the date of retirement. The plan is unfunded.

Young Ma was hired at the beginning of 1995 and is expected to retire at the end of 2023. The discount rate is 4%.

The PV of an ordinary annuity of $1 where n = 15 and i = 4% is 11.11839.

The PV of $1 where n = 2 and i = 4% is 0.92456.

With respect to Ma, what is the service cost in ABC’s 2021 postretirement benefit expense? (Hint: First calculate EPBO, then discount EPBO to 2021, then divide by the years in the attribution period).

A) $7,089.

B) $6,365.

C) $23,000.

D) $5,272.
Answer #1

Annuity Amount = 20000 annually for 15 years

PV at the end of 2023 (retirement date)= 20000 * 11.11839= 222367.8 (EPBO)

PV at the end of 2021= 222367.8 * 0.92456 =205592.37

Attribution period= 29 years (1995 to 2023)

Service cost in ABC's 2021 postretirement benefit expense = 205592.37 / 29 = 7089

Answer: A) 7089